OpenStudy (anonymous):

2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

just add like terms together to simplify this expression >.< so like terms would be; 2x, -2x -2y, -y 5z, 3z so 2x - 2x = 0 -2y - y = -3y 5z + 3z = 8z so then there's 0 -3y + 8z -3y + 8z

OpenStudy (marissalovescats):

Welcome to OpenStudy @simon182__ !